OpenStudy (anonymous):

Solve: 4(x + 1) = −12 −2 4 −3.25 −4

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

grr is cool

OpenStudy (zzr0ck3r):

4(X+1) = -12 4X+4 = -12 4X = -16 X = -16/4 X=?

OpenStudy (zzr0ck3r):

are we sapposed to just give the answer?

OpenStudy (anonymous):

Zzr0cker how did you get 4X +4 = -12 ?

OpenStudy (anonymous):

wouldnt it be 8x ?

OpenStudy (zzr0ck3r):

distribute the 4 threw

OpenStudy (zzr0ck3r):

a(x+b) = ax+ab

OpenStudy (anonymous):

Yes i know that 4 times x = 4x and 4 times 1 = 4 but whats next ? this is where im stuck

OpenStudy (anonymous):

\[\small4(x + 1) = -12 \implies 4(x) + 4(1) = -12 \implies 4x + 4 = -12\implies 4x = -16 \implies x = -4\]\[\Huge OR\]\[4(x + 1) = -12 \implies x + 1 = -3 \implies x = -4\]

OpenStudy (anonymous):

wouldnt 4x + 4 = 8x ? i dont get it ...

OpenStudy (anonymous):

Where are you getting the 8x from?

OpenStudy (anonymous):

Cause if you had 4x + 4 wouldnt that = 8x ?

OpenStudy (anonymous):

No. 4x and 4 are not like terms since 4 does not have an x. Plus, there is a -12 on the other side.

OpenStudy (anonymous):

It would onyl equal 8x if it were 4x + 4x.

OpenStudy (zzr0ck3r):

4 apples + 4 apples = 8 apples 4apples + 4 does not equal 8 apples

OpenStudy (zzr0ck3r):

its sort of odd because of the way we speak. Like if i say I have 4 apples and someone says add 4. You assume they mean add 4 apples. but with math we cant assume:)